what is 66 2/3 percent of 27
Answer by KnightOwlTutor(293)

Convert 2/3 into decimal format =0.667
Add 66+.0667=66.667
To convert the percentage into decimal =66.667/100=0.6667=0.67
Multiply (0.67)(27)=18.09